Q1. Determine the inverse Laplace Transform a) I(s) = 2125+5 9(5+1) (5F

c) V() = x2 +592 +28 5152455+2) 65+ 24 = 65+241
Answer #1

30 Saluhiona) IC) = g(+1) 9 $+1) (82423+5 (2+1)²42² I(S) = q (+1) (S +1 +21) (5+1-21) IC) = 0.05+13 Şts +1 +2139693-223 By pFC) = 2x8+(47)3 + + $5+(-2013 10)=2576172+ 5+ (0-20) taking inves se tap of transform. Laplace (C)= q e64203€ 01-27)+7 2 RectSeluhan: 65424) (5² 453 ²4 25 W) = 65+4) 205245542) lot orta) - 66ta) SiS+0.438)(3+4.556) By partial tractions VG) = + ( 38 4VC)-$12, -12 0.439) 0-2 ? (5+4.55) S VO)= 12 12 0-2 I (a to 830) (2+257) taking Inverse lablace foams from Ut)= 12-12 0043&t.